# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ Abstract
Despite linear regression being the most popular statistical modelling technique, in real-life we often need to deal with situations where the true relationship between the response and the covariates is nonlinear in parameters. In such cases one needs to adopt appropriate non-linear regression (NLR) analysis, having wider applications in biochemical and medical studies among many others. In this paper we propose a new improved robust estimation and testing methodologies for general NLR models based on the minimum density power divergence approach and apply our proposal to analyze the widely popular Michaelis-Menten (MM) model in enzyme kinetics. We establish the asymptotic properties of our proposed estimator and tests, along with their theoretical robustness characteristics through influence function analysis. For the particular MM model, we have further empirically justified the robustness and the efficiency of our proposed estimator and the testing procedure through extensive simulation studies and several interesting real data examples of enzyme-catalyzed (biochemical) reactions.
@@ 1 Introduction
Linear regression is arguably the most used statistical modeling technique to investigate relationships between variables. However, even for normally distributed continuous response variables, the linearity assumption appears to be restrictive in several real-life situations where the true relationship between the response and the covariates may be best modelled by a suitable non-linear function in parameters. Such scenarios frequently arise in biomedical and biochemical researches, besides statistical signal analyses and many other domains. For example, a popular such non-linear relationship in enzyme kinetics is known as the Michaelis-Menten (MM) equation. This MM equation, originally developed by Leonor Michaelis and Maud Menten in 1913 [30], has now become the foundational basis of enzymology for studying the rates of enzyme catalyzed chemical reactions; see, e.g., [2, 3, 22, 31, 35]. For such a reaction, if the reaction velocity is $v$ at a substrate concentration $s$, then the MM equation says
@@
where $K_{m}$ is the pseudo-equilibrium constant of the reaction, called Michaelis constant, and $V_{max}$ is the maximum velocity attainable (theoretically) at an infinite substrate concentration saturated with the enzyme. Here $K_{m}$ and $V_{max}$ are the two reaction-specific unknown (model) parameters which are estimated from experimental data, consisting of the observed reaction velocities ($v_{i}$) at different (often pre-fixed) substrate concentrations ($s_{i}$), for a specific chemical reaction, assuming additive (and independent) Gaussian error in each observation $v_{i}$ (for $i=1,\ldots,n$, say). This estimation setup can be expressed as a non-linear regression (NLR) model given by
$v_{i}=\mu(s_{i},\bm{\beta})+\epsilon_{i},~{}~{}~{}~{}i=1,\ldots,n,$ (2)
where $\mu(s_{i},\bm{\beta})=\frac{\beta_{1}s_{i}}{\beta_{2}+s_{i}}$ and $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$. The estimation of the model parameter $\bm{\beta}$ and the subsequent inference are commonly performed using the non-linear least-squares or the maximum likelihood approach [18, 37, 21]. These inferences are also extended to variants of the MM model for specific applications; see, e.g., [38, 12]. The MM model (1) can also be analyzed using a linearizing transformation, called Woolf transformation, used by Currie [11]. However, for that case it is not possible to take additive errors, as in model (2).
More generally, given $n$ observations $y_{i}$ and $\bm{x}_{i}$ on a (real valued) response variable $Y$ and associated $k$-dimensional covariates $\bm{X}$, respectively, for $i=1,\ldots,n$, a typical NLR model can be expressed as [9, 34]
@@
where $\bm{\beta}$ is a $p$-dimensional regression parameter vector and $\epsilon_{i}$’s are independently and identically distributed (IID) random errors with mean 00 and variance $\sigma^{2}$. Here the form of the mean (or regression) function $\mu(\cdot,\cdot)$ is assumed to be known (pre-specified), continuous, and non-linear in the parameter $\bm{\beta}$. The full parameter vector in this NLR model is then given by $\bm{\theta}=(\bm{\beta}^{T},\sigma^{2})^{T}\in\bm{\Theta}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{+}$. It is to be noted that, since $\mu(\bm{x_{i}},\bm{\beta})$ is non-linear in $\bm{\beta}$, dimensions of $\bm{x_{i}}$ and $\bm{\beta}$ are not necessarily the same (i.e., $p$ may or may not be equal to $k$). There is a vast amount of literature on the classical ordinary least-square (OLS) and the maximum likelihood (ML) methods under NLR models. Asymptotic properties of the least-square estimator under the NLR model were initially studied by Jennrich [20] and Wu [40]. It can also be easily observed that, under the normality assumption on the IID errors $\epsilon_{i}$’s in (3), both the OLS estimator and the ML estimator (MLE) of the parameter $\bm{\beta}$ coincide. Discussions on the MLE under the NLR model (3) and its asymptotic properties are available in, e.g., Seber and Wild [34]. Further refinements and advancements in both theory and computation of these estimators under general NLR models and several of its important special cases are available in, e.g., [32, 23, 24, 25], among many others.
@@
However, the main drawback of the least-squares or the maximum likelihood based inference is their high sensitivity to the presence of outliers and other contaminations in the observed data, even in a small proportion. Since we frequently encounter outliers and influential observations in real-life datasets, there is a clear need for suitable robust estimation and inference procedures yielding stable and accurate results even under data contamination. But, although there exists a vast literature on robust statistical procedures for linear and generalized linear regression models, there have been comparatively fewer works developing robust procedures for NLR models. In particular, Liu et al. [27] and Marasovic et al. [28] studied robust M-estimators under the NLR models with Huber’s and Tukey’s loss/weight functions, respectively. Tabatabai et al. [36] proposed a modified M-estimation procedure (to be referred to as the KPS estimator) whereas Liu et al. [26] studied the median of means (MOM) estimator under the NLR model. Giradi et al. [17] used Tsallis score for robust estimation under a particular example of 3-parameter log-logistic NLR model, for analyzing COVID-19 contagion data of Italy, but they avoided many mathematical details and justifications of the estimation procedure. Further, all these existing works only dealt with the estimation problem and there is no concrete work on robust hypotheses testing available in the literature of NLR models. In the context of MM models in enzyme kinetics, only the KPS and the M-estimators are explored to generate robust solutions under possible data contaminations; see, e.g., [36, 28, 27], among others.
@@
In this paper our aim is to develop an improved robust estimator and associated robust testing procedure for the general class of NLR models based on the popular minimum density power divergence (DPD) approach of Basu et al. [6], and apply them for robust analysis of MM models with greater accuracy compared to the existing alternatives. Due to various nice properties, the DPD based approach has been successfully implemented in several statistical problems, including regressions; see, e.g., [14, 15, 16, 7, 5, 13, 10, 4]. For the sake of completeness, a brief background about the DPD and the minimum DPD estimator (MDPDE) is provided in Appendix A.1. Here we will extend the definition of the MDPDE for the general NLR models and discuss its asymptotic properties (consistency and asymptotic normality) under simple verifiable conditions. We will justify the robustness property of our proposed estimator by examining the boundedness of its influence function. We will additionally develop a robust hypothesis testing procedure for general NLR models based on our proposed estimator (i.e., the MDPDE) and examine the influence functions of our test statistics to justify their claimed robustness. We will then apply our proposed methodologies for robust inference under the MM model (2), with simplified theoretical properties and extensive empirical illustrations of their finite-sample performances. In particular, we will develop and study a robust test procedure for testing the validity of the MM model for a given (possibly contaminated) experimental dataset. Finally, we will present some interesting real data examples of different enzyme kinetic reactions.
@@
The rest of the paper is organized as follows: In Section 2, we present our proposed MDPDE along with the derivation of its asymptotic and robustness properties. In Section 3 we have discussed the proposed robust hypothesis testing procedures utilizing the MDPDEs and studied their asymptotic properties and their influence functions. Section 4 is devoted to the application of our proposed methods to the Michaelis-Menten model (2) along with extensive simulation studies illustrating their improved finite-sample performances in comparison to the existing robust approaches. Section 5 consists of discussion on some data driven procedures for finding an optimal choice of the MDPDE tuning parameter. Section 6 presents analyses of some real data examples and final concluding remarks are given in Section 7. For brevity in presentation, some background details, proofs of the results, and some numerical illustrations are provided in Appendices A-C.
@@ 2.1 The MDPDE
Let us consider the NLR model given in (3) with the error assumption $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently, for all $i=1,2,\ldots,n$. So, we have $n$ independent and non-homogeneous (INH) observations $y_{1},y_{2},\ldots,y_{n}$ with
$y_{i}\sim\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right),\mbox{ independently, for }i=1,\ldots,n.$
So, we can follow the general theory of the MDPDE under INH setup, as described briefly in Appendix A.1 following Ghosh and Basu [14], to define and study the MDPDE of the parameters under the present NLR model. Continuing with the notations from Appendix A.1, in this case, we have the model density for $i$-th observation as $f_{i}(\cdot;\bm{\theta})\equiv\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right)$, so that we get
@@
Now, differentiating (5) partially with respect to $\bm{\beta}$ and $\sigma^{2}$, and equating to 00, we get the estimating equations for the MDPDE as given by
@@
where $\nabla_{\bm{\beta}}$ denotes the first order partial derivative with respect to $\bm{\beta}$. To obtain the MDPDE $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ with tuning parameter $\alpha>0$, in practice, one can numerically solve the above two estimating equations simultaneously for $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$, which is equivalent to minimize (5). Note that, the MDPDE coincides with the MLE at $\alpha=0$ in a limiting sense, and provides a robust generalization of the MLE at $\alpha>0$.
Note that, an important advantage of our MDPDE over its existing robust competitors (briefly described in Appendix A.2) is that the MDPDEs additionally yield a robust estimate of the scale parameter $\sigma$ simultaneously with the estimation of the regression parameter $\bm{\beta}$. Further, for the MDPDE, the problem with multiple roots of the estimation equations can be easily solved using the associated objective function in (5).
@@ 2.2 Asymptotic properties
For simplicity, let us assume that the true data generating density $g_{i}$ of $y_{i}$ belongs to the corresponding parametric model family
$\mathcal{F}_{i}=\left\{\mathcal{N}\left(\mu(\bm{x_{i}},\bm{\beta}),\sigma^{2}\right):\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}\in\bm{\Theta}\right\},~{}~{}~{}i=1,2,\ldots,n,$ (8)
i.e., $g_{i}\equiv\mathcal{N}\left(\mu(\bm{x_{i},\beta_{0}}),\sigma_{0}^{2}\right),\mbox{ for some }\bm{\theta_{0}}=\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}$. Our NLR model (3) can be expressed in vectorized form as
$\bm{y}=\bm{\mu(\beta)}+\bm{\epsilon},$ (9)
where $\bm{y}=(y_{1},\ldots,y_{n})^{T},\bm{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ and $\bm{\mu(\beta)}=\left(\mu(\bm{x_{1},\beta}),\ldots,\mu(\bm{x_{n},\beta})\right)^{T}$. Now we define the $n\times p$ matrix
$\bm{\dot{\mu}(\beta)}=\frac{\partial}{\partial\bm{\beta}}\bm{\mu(\beta)}=\begin{pmatrix}\nabla_{\bm{\beta}}\mu(\bm{x_{1}},\bm{\beta})^{T}\\ \nabla_{\bm{\beta}}\mu(\bm{x_{2}},\bm{\beta})^{T}\\ \vdots\\ \nabla_{\bm{\beta}}\mu(\bm{x_{n}},\bm{\beta})^{T}\\ \end{pmatrix},$ (10)
and the $(p+1)\times(p+1)$ matrices
$\bm{\Psi}_{n}=\begin{bmatrix}\frac{\zeta_{\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{\alpha}\end{bmatrix}\mbox{ and }\bm{\Omega}_{n}=\begin{bmatrix}\frac{\zeta_{2\alpha}}{n}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)&\bm{0}_{p}\\ \\ \bm{0}_{p}^{T}&\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}\end{bmatrix},$ (11)
where
$\zeta_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+2)}(1+\alpha)^{-3/2}\mbox{ and }\varsigma_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+4)}\frac{2+\alpha^{2}}{4(1+\alpha)^{5/2}}.$
We now present some conditions on the NLR model, under which the asymptotic properties of the MDPDE $\bm{\widehat{\theta}_{\alpha}}$ would be derived. Let us denote $\bm{\Theta_{\beta}}\ (\subseteq\mathbb{R}^{p})$ be the parameter space of the model parameter $\bm{\beta}$ of the NLR model (3) and $\sigma^{2}>0$, so that $\bm{\Theta}=\bm{\Theta_{\beta}}\times\mathbb{R}^{+}$.
@@
(R1) Let $(\bm{\dot{\mu}})_{ij}$ be the $(i,j)$-th element of the matrix $\bm{\dot{\mu}(\beta)}$, given in (10). Then, for all $j,k,l$ and for all $\bm{\beta}\in\omega$, an open subset of $\bm{\Theta_{\beta}}$, the mean function $\mu(\bm{x_{i}},\bm{\beta})$ is thrice continuously differentiable with respect to $\bm{\beta}$ with
$\sup_{n>1}\max_{1\leq i\leq n}\left|(\bm{\dot{\mu}})_{ij}\right|=O(1),\ \ \ \ \sup_{n>1}\max_{1\leq i\leq n}\left|(\bm{\dot{\mu}})_{ij}(\bm{\dot{\mu}})_{ik}\right|=O(1),$ (12)
and
$\frac{1}{n}\sum_{i=1}^{n}\left|(\bm{\dot{\mu}})_{ij}(\bm{\dot{\mu}})_{ik}(\bm{\dot{\mu}})_{il}\right|=O(1).$ (13)
(R2) The matrix $\bm{\dot{\mu}(\beta)}$ satisfies
$\inf_{n}\left[\mbox{minimum eigenvalue of }\frac{1}{n}\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right]>0,$ (14)
which says that $\bm{\dot{\mu}(\beta)}$ is of full column rank, and
$n\max_{1\leq i\leq n}\left[\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})^{T}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})\right]=O(1).$ (15)
With the above mentioned conditions, the asymptotic distributions of the MDPDEs of the parameters of the NLR model (3) are given in the following theorem; its proof is given in Appendix B.1.
@@
Under the setup of the NLR model in (3), suppose that the true data generating distributions belong to the corresponding parametric model families with the true parameter value being $\bm{\theta}_{0}=\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}$ and Conditions (R1)-(R2) hold. Then, for any fixed $\alpha\geq 0$, we have the following results.
@@
    There exists a consistent sequence of roots $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ of the MDPDE estimating equations (6) and (7).

### v2

@@ Abstract
Despite linear regression being the most popular statistical modelling technique, in real-life we often need to deal with situations where the true relationship between the response and the covariates is nonlinear in parameters. In such cases, one needs to adopt appropriate non-linear regression (NLR) analysis, having wider applications in biochemical and medical studies among many others. In this paper, we propose a new improved robust estimation and testing methodologies for general NLR models based on the minimum density power divergence approach and apply our proposal to analyze the widely popular Michaelis-Menten (MM) model in enzyme kinetics. We establish the asymptotic properties of our proposed estimator and tests, along with their theoretical robustness characteristics through influence function analysis. For the particular MM model, we have further empirically justified the robustness and the efficiency of our proposed estimator and the testing procedure through extensive simulation studies and several interesting real data examples of enzyme-catalyzed (biochemical) reactions.
@@ 1 Introduction
Linear regression is arguably the most used statistical modeling technique to investigate relationships between variables. However, even for normally distributed continuous response variables, the linearity assumption appears to be restrictive in several real-life situations where the true relationship between the response and the covariates may be best modeled by a suitable non-linear function in parameters. Such scenarios frequently arise in biomedical and biochemical researches, besides statistical signal analyses and many other domains. For example, a popular such non-linear relationship in enzyme kinetics is known as the Michaelis-Menten (MM) equation. This MM equation, originally developed by Leonor Michaelis and Maud Menten in 1913 [33], has now become the foundational basis of enzymology for studying the rates of enzyme-catalyzed chemical reactions; see, e.g., [2, 3, 24, 34, 38]. For such a reaction, if the reaction velocity is $v$ at a substrate concentration $s$, then the MM equation says
@@
where $K_{m}$ is the pseudo-equilibrium constant of the reaction, called the Michaelis constant, and $V_{max}$ is the maximum velocity attainable (theoretically) at an infinite substrate concentration saturated with the enzyme. Also, from Eq. (1), we observe that $v=V_{max}/2$ if and only if $K_{m}=s$, and hence the Michaelis constant $K_{m}$ can further be interpreted as the substrate concentration needed to attain half of the maximum attainable velocity $V_{max}$. Here $K_{m}$ and $V_{max}$ are the two reaction-specific unknown (model) parameters that are estimated from experimental data, consisting of the observed reaction velocities ($v_{i}$) at different (often pre-fixed) substrate concentrations ($s_{i}$), for a specific chemical reaction, assuming additive (and independent) Gaussian error in each observation $v_{i}$ (for $i=1,\ldots,n$, say). This estimation setup can be expressed as a non-linear regression (NLR) model given by
$v_{i}=\frac{\beta_{1}s_{i}}{\beta_{2}+s_{i}}+\epsilon_{i}=\mu(s_{i},\bm{\beta})+\epsilon_{i},~~~~~~~i=1,\ldots,n,$ (2)
where $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$. The estimation of the model parameter $\bm{\beta}$ and the subsequent inference is commonly performed using the non-linear least-squares or the maximum likelihood approach; see e.g., [19, 40, 23]. These inferences are also extended to variants of the MM model for specific applications; see, e.g., [41, 12]. The MM model (1) can also be analyzed using a linearizing transformation, called the Woolf transformation, used by Currie [11]. However, in that case, it is not possible to take additive errors, as in model (2).
More generally, given $n$ observations $y_{i}$ and $\bm{x}_{i}$ on a (real-valued) response variable $Y$ and associated $k$-dimensional covariates $\bm{X}$, respectively, for $i=1,\ldots,n$, a typical NLR model can be expressed as [9, 37]
@@
where $\bm{\beta}$ is a $p$-dimensional regression parameter vector and $\epsilon_{i}$’s are independently and identically distributed (IID) random errors with mean 0 and variance $\sigma^{2}$. Here the form of the mean (or regression) function $\mu(\cdot,\cdot)$ is assumed to be known (pre-specified), continuous, and non-linear in the parameter $\bm{\beta}$. The full parameter vector in this NLR model is then given by $\bm{\theta}=(\bm{\beta}^{T},\sigma^{2})^{T}\in\bm{\Theta}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{+}$. It is to be noted that, since $\mu(\bm{x}_{i},\bm{\beta})$ is non-linear in $\bm{\beta}$, dimensions of $\bm{x}_{i}$ and $\bm{\beta}$ are not necessarily the same (i.e., $p$ may or may not be equal to $k$). There is a vast amount of literature on the classical ordinary least-square (OLS) and the maximum likelihood (ML) methods under NLR models. Asymptotic properties of the least-square estimator under the NLR model were initially studied by Jennrich [22] and Wu [43]. It can also be easily observed that under the normality assumption on the IID errors $\epsilon_{i}$’s in (3), both the OLS estimator and the ML estimator (MLE) of the parameter $\bm{\beta}$ coincide. Discussions on the MLE under the NLR model (3) and its asymptotic properties are available in, e.g., Seber and Wild [37]. Further refinements and advancements in both theory and computation of these estimators under general NLR models and several of its important special cases are available in, e.g., [35, 26, 27, 25], among many others.
@@
However, the main drawback of the least-squares or the maximum likelihood-based inference is their high sensitivity to the presence of outliers and other contaminations in the observed data, even in a small proportion. Since we frequently encounter outliers and influential observations in real-life datasets, there is a clear need for suitable robust estimation and inference procedures yielding stable and accurate results even under data contamination. But, although there exists a vast literature on robust statistical procedures for linear and generalized linear regression models, there have been comparatively fewer works developing robust procedures for NLR models. In particular, Liu et al. [29] and Marasovic et al. [31] studied robust M-estimators under the NLR models with Huber’s and Tukey’s loss/weight functions, respectively. Tabatabai et al. [39] proposed a modified M-estimation procedure (to be referred to as the KPS estimator), whereas Liu et al. [28] studied the median of means (MOM) estimator under the NLR model. Girardi et al. [18] used the Tsallis score for robust estimation under a particular example of 3-parameter log-logistic NLR model, for analyzing COVID-19 contagion data of Italy, but they avoided many mathematical details and justifications of the estimation procedure. Further, all these existing works only dealt with the estimation problem, and there is no concrete work on robust hypotheses testing available in the literature of NLR models. In the context of MM models in enzyme kinetics, only the KPS and the M-estimators are explored to generate robust solutions under possible data contaminations; see, e.g., [39, 31, 29], among others.
@@
In this paper, our aim is to develop an improved robust estimator and associated robust testing procedure for the general class of NLR models based on the popular minimum density power divergence (DPD) approach of Basu et al. [6], and apply them for robust analysis of MM models with greater accuracy compared to the existing alternatives. Due to various nice properties, the DPD-based approach has been successfully implemented in several statistical problems, including regressions; see, e.g., [13, 14, 15, 7, 5, 17, 10, 4]. For the sake of completeness, a brief background about the DPD and the minimum DPD estimator (MDPDE) is provided in the online supplementary material. Here, we extend the definition of the MDPDE for the general NLR models and discuss its asymptotic properties (consistency and asymptotic normality) under simple verifiable conditions. We justify the robustness property of our proposed estimator by examining the boundedness of its influence function. We additionally develop a robust hypothesis testing procedure for general NLR models based on our proposed estimator (i.e., the MDPDE) and examine the influence functions of our test statistics to justify their claimed robustness. We then apply our proposed methodologies for robust inference under the MM model (2), examine their theoretical properties by further simplifying the formulae for asymptotic variances and influence functions, and illustrate their finite-sample performances through extensive simulation studies. We also develop and study a robust test procedure for testing the validity of the MM model for a given (possibly contaminated) experimental dataset. Further, we discuss some data-driven procedures for finding an optimal choice of the MDPDE tuning parameter in the online supplementary material. Finally, we present some interesting real data examples of different enzyme kinetic reactions.
@@
The rest of the paper is organized as follows: In Section 2, we present our proposed MDPDE along with the derivation of its asymptotic and robustness properties. In Section 3, we discuss the proposed robust hypothesis testing procedures utilizing the MDPDEs and study their asymptotic properties and their influence functions. Section 4 is devoted to the application of our proposed methods to the Michaelis-Menten model (2) along with extensive simulation studies illustrating their improved finite-sample performances in comparison to the existing robust approaches. Section 5 consists of a discussion on some data-driven procedures for finding an optimal choice of the MDPDE tuning parameter. Section 6 presents analyses of some real data examples, and final concluding remarks are given in Section 7. For brevity in the presentation, some background details, proofs of the results, and some numerical illustrations are provided in the web appendices of the online supplementary material.
@@ 2.1 The MDPDE
Let us consider the NLR model given in (3) with the error assumption $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently, for all $i=1,2,\ldots,n$. So, we have $n$ independent and non-homogeneous (INH) observations $y_{1},y_{2},\ldots,y_{n}$ with $y_{i}\sim\mathcal{N}\left(\mu(\bm{x}_{i},\bm{\beta}),\sigma^{2}\right)$, independently, for $i=1,\ldots,n$. So, we can follow the general theory of the MDPDE under INH setup, as described briefly in Web Appendix A.1, to define and study the MDPDE of the parameters under the present NLR model. Following the notations of Web Appendix A.1, the model density for $i$-th observation is $f_{i}(\cdot;\bm{\theta})\equiv\mathcal{N}\left(\mu(\bm{x}_{i},\bm{\beta}),\sigma^{2}\right)$ and we have
@@
Now, differentiating (5) partially with respect to $\bm{\beta}$ and $\sigma^{2}$, and equating to 0, the estimating equations for the MDPDE are obtained as
@@
where $\nabla_{\bm{\beta}}$ denotes the first order partial derivative with respect to $\bm{\beta}$. To obtain the MDPDE $\bm{\widehat{\theta}}_{\alpha}=(\bm{\widehat{\beta}}_{\alpha}^{T},\widehat{\sigma}_{\alpha}^{2})^{T}$ with tuning parameter $\alpha>0$, in practice, one can numerically solve the above two estimating equations simultaneously for $\bm{\theta}=(\bm{\beta}^{T},\sigma^{2})^{T}$, which is equivalent to minimize (5). Clearly, the MDPDE coincides with the MLE at $\alpha=0$ in a limiting sense, and provides a robust generalization of the MLE at $\alpha>0$.
Note that, an important advantage of our MDPDE over its existing robust competitors (briefly described in Web Appendix A.2) is that the MDPDEs additionally yield a robust estimate of the scale parameter $\sigma$ simultaneously with the estimation of the regression parameter $\bm{\beta}$. Further, unlike other estimating equation based estimators, the MDPDE is associated with a nice objective function, which helps us to avoid the problem of multiple roots of the estimating equations. If we get multiple solutions to the MDPDE estimating equations (6)-(7), we should choose the one yielding the minimum value of the associated objective function given in (5). The same can also be used while choosing between multiple local minima of the said objective function.
@@ 2.2 Asymptotic properties
For simplicity, let us assume that the true data-generating density $g_{i}$ of $y_{i}$ belongs to the corresponding parametric model family $\mathcal{F}_{i}=\{\mathcal{N}(\mu(\bm{x}_{i},\bm{\beta}),\sigma^{2}):\bm{\theta}=(\bm{\beta}^{T},\sigma^{2})^{T}\in\bm{\Theta}\}$, $i=1,2,\ldots,n,$ i.e., $g_{i}\equiv\mathcal{N}\left(\mu(\bm{x}_{i},\bm{\beta}_{0}),\sigma_{0}^{2}\right),\mbox{ for some }\bm{\theta}_{0}=\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}$. Our NLR model (3) can be expressed in vectorized form as
$\bm{y}=\bm{\mu}(\bm{\beta})+\bm{\epsilon},$ (8)
where $\bm{y}=(y_{1},\ldots,y_{n})^{T},\bm{\epsilon}=(\epsilon_{1},\ldots,\epsilon_{n})^{T}$ and $\bm{\mu}(\bm{\beta})=\left(\mu(\bm{x}_{1},\bm{\beta}),\ldots,\mu(\bm{x}_{n},\bm{\beta})\right)^{T}$. Now we define the $n\times p$ matrix
$\bm{\dot{\mu}}(\bm{\beta})=\frac{\partial\bm{\mu}(\bm{\beta})}{\partial\bm{\beta}^{T}}=\begin{bmatrix}\nabla_{\bm{\beta}}\mu(\bm{x}_{1},\bm{\beta})&\nabla_{\bm{\beta}}\mu(\bm{x}_{2},\bm{\beta})&\ldots&\nabla_{\bm{\beta}}\mu(\bm{x}_{n},\bm{\beta})\\ \end{bmatrix}^{T},$ (9)
where $\nabla_{\bm{\beta}}$ denotes the partial derivative with respect to $\bm{\beta}$, and the $(p+1)\times(p+1)$ matrices
$\bm{\Psi}_{n}=\begin{bmatrix}\frac{\zeta_{\alpha}}{n}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)&\bm{0}_{p}\\ \bm{0}_{p}^{T}&\varsigma_{\alpha}\end{bmatrix}\mbox{ and }\bm{\Omega}_{n}=\begin{bmatrix}\frac{\zeta_{2\alpha}}{n}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)&\bm{0}_{p}\\ \bm{0}_{p}^{T}&\varsigma_{2\alpha}-\frac{\alpha^{2}}{4}\zeta_{\alpha}^{2}\end{bmatrix}$, (10)
where $\zeta_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+2)}(1+\alpha)^{-3/2}$ and $\varsigma_{\alpha}=(2\pi)^{-\alpha/2}\sigma^{-(\alpha+4)}\frac{2+\alpha^{2}}{4(1+\alpha)^{5/2}}$. Let us further denote the $(i,j)$-the entry of the matrix $\bm{\dot{\mu}}(\bm{\beta})$ by $\dot{\mu}_{ij}(\bm{\beta})$ for all $i=1,\ldots n$ and $j=1,\ldots,p$.
We now present some conditions on the NLR model, under which the asymptotic properties of the MDPDE $\bm{\widehat{\theta}}_{\alpha}$ would be derived. Let us denote $\bm{\Theta_{\beta}}\ (\subseteq\mathbb{R}^{p})$ be the parameter space of the model parameter $\bm{\beta}$ of the NLR model (3) and $\sigma^{2}>0$. So, $\bm{\Theta}=\bm{\Theta_{\beta}}\times\mathbb{R}^{+}$.
@@
The mean function $\mu(\bm{x}_{i},\bm{\beta})$ is thrice continuously differentiable with respect to $\bm{\beta}\in\omega$, an open subset of $\bm{\Theta_{\beta}}$ containing the true parameter value ($\bm{\beta}_{0}$). Further, for all $i,j,k,l$ and $\bm{\beta}\in\omega$, we have
$\sup_{n>1}\max_{1\leq i\leq n}\left|\dot{\mu}_{ij}(\bm{\beta})\right|=O(1),\ \ \ \ \sup_{n>1}\max_{1\leq i\leq n}\left|\dot{\mu}_{ij}(\bm{\beta})\dot{\mu}_{ik}(\bm{\beta})\right|=O(1),$ (11)
$\mbox{and}~~~~~~~~~~\frac{1}{n}\sum_{i=1}^{n}\left|\dot{\mu}_{ij}(\bm{\beta})\dot{\mu}_{ik}(\bm{\beta})\dot{\mu}_{il}(\bm{\beta})\right|=O(1).$~~~~~~~~~~ (12)
(R2) For all $\bm{\beta}\in\omega$, the matrix $\bm{\dot{\mu}}(\bm{\beta})$ satisfies
$\inf_{n}\left[\mbox{minimum eigenvalue of }\frac{1}{n}\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right]>0,$ (13)
which says that $\bm{\dot{\mu}}(\bm{\beta})$ is of full column rank, and
$\max_{1\leq i\leq n}\left[\nabla_{\bm{\beta}}\mu(\bm{x}_{i},\bm{\beta})^{T}\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{-1}\nabla_{\bm{\beta}}\mu(\bm{x}_{i},\bm{\beta})\right]=O(n^{-1}).$ (14)
With the above-mentioned conditions, the asymptotic distributions of the MDPDEs of the NLR model parameters are given in the following theorem; its proof is given in Web Appendix B.
@@
Under the setup of the NLR model (3), suppose that the true data-generating distributions belong to the corresponding parametric model families with the true parameter value being $\bm{\theta}_{0}=\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}$ and Conditions (R1)-(R2) hold. Then, for any $\alpha\geq 0$, we have the following results.
@@
    There exists a consistent sequence of roots $\bm{\widehat{\theta}}_{\alpha}=\left(\bm{\widehat{\beta}}_{\alpha}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ of the MDPDE estimating equations (6)–(7).